You must decide whether to buy a new car for $20,000 or less the same car over a three-year. Under the terms of the lease you can down make a down payment of $3,000 and have monthly payments of $200 at the end of the three years the lease car has a residual value that's the amount you pay if you choose to buy the car at the end of the lease. Of $12,000 assume you can sell the new car at the end of the three years at the same residual value what is the cost for buying the car and selling it after 3 years.$

Answer by Theo(13342)

given the figures you provided, and assuming that you will sell the car at the end of the 3 year period, and assuming that the financing on the car is for 3 years 5% per year compounded monthly, and assuming you make the same down payment on buying the car as on leasing the car, and assuming that you can sell the car for the residual value of 12,000 at the end of the buy period, and assuming no other costs are indicted in either buy or lease consideration, then my analysis shows the following:

lease option.

down payment 3000
monthly payments 200 a month for 36 months.
total cost = 10,200.

buy option.

down payment 3000
loan amount = 20,000 - 3,000 = 17,000
interest rate at 2%, 3%, 4%, 5%, 6%.
value of car after 3 years = 12,000.

net cash expenses, other than the loan, are + 3,000 down payment minus 12,000 for sale of car after 3 years of ownership = -9,000.

total cost at different interest rates on the loan.

2% = 486.92 * 36 - 9,000 = 8,529.12
3% = 494.38 * 36 - 9,000 = 8,797.68
4% = 501.92 * 36 - 9,000 = 9,069.12
5% = 509.51 * 36 - 9,000 = 9,345.96
6% = 517.17 * 36 - 9,000 = 9,618.12

based on your figures, and assuming no additional cost differences than the ones presented, it appears the lease versus buy cost comparison favors the buy option.

a complete analysis is not that simple, but this give you a rough idea.